OpenStudy (anonymous):

Factor completely:8x^3-125 all steps please!

OpenStudy (anonymous):

Use the sum of two cubes: 8x^3+125 = (2x)^3 + 5^3 then, u get...

OpenStudy (anonymous):

(2x+5)(4x^2-10x+25).. and thats how u do it:)

OpenStudy (ajprincess):

\(8x^3-125\) \(=2^3x^3-5^3\) \(=(2x)^3-5^3\) Difference of cubes. \(a^3-b^3=(a-b)(a^2+ab+b^2)\) Here a=2x and b=5